OpenStudy (anonymous):

write an equation in the point-slope form and the slope-intercept form. slope=4, passing through (1,3)

OpenStudy (mew55):

using the formula y-y0=m(x-x0) y-3=4(x-1) y-3+3=4x-4+3 y=4x-1

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

write an equation in the point-slope form and the slope-intercept form. slope=8, passing through (4,-1) slope=-5, passing through (-4,-2) slope-2, passing through (0, -3)